A 20 Newton force is exerted for 2 seconds on an object that initially has 50 kg * m/s of momentum. How much momentum does the object have after the force is exerted?


1
Expert's answer
2021-04-26T17:11:36-0400
"p_f=p_i+Ft\\p_f=50+20(2)=90\frac{kgm}{s}"